OpenStudy (igster1112):

Step by step. he do you write a rule to describe a transformation? Such as 1 up and 1 right.....or 3 down and 1 right

OpenStudy (amistre64):

http://openstudy.com/study#/updates/540b4d79e4b081860be4fb79 i spose i was taking to long?

OpenStudy (igster1112):

Nah it's fine

OpenStudy (amistre64):

up and down are y movements left and right are x movements just add/subtract as needed

OpenStudy (igster1112):

So how do i tell if the numbers are positive or negative in the coordinate?

OpenStudy (amistre64):

lets say T maps the point (x,y) to a new point (f(x),f(y)) such that we go up by 4 and left by 3 then f(x) = x-3 and f(y) = y+4

OpenStudy (amistre64):

in general, up and right are positive directions down and left are negative directions

OpenStudy (amistre64):

if you can draw a 'graph' the y axis is up and down, the x axis is left and right, and they cross at the zeros 

OpenStudy (igster1112):

Oh ok

OpenStudy (igster1112):

Had go find the rule for 1 up and 1 right and i put (1 -1) and the teacher put (-1 1)....I need a practice one.

OpenStudy (amistre64):

yeah, x first; the y

OpenStudy (amistre64):

but 1 up and 1 right are both postive directions so even if you swapped them they are the same

OpenStudy (amistre64):

T:(x,y) to (x+1,y+1)

OpenStudy (igster1112):

Oh ok

OpenStudy (igster1112):

So I would put the left and right and then the Up and down?

OpenStudy (igster1112):

So if i was doing 3 down and one right which number would I put first?

OpenStudy (igster1112):

I got it...it would be (1-3)?